HelpMePlz:

At the start of January, you have a credit card with a balance of $498, a credit limit of $900, and an interest rate of 15%. You make monthly payments of $35. At the end of which month will your balance be below the target debt ratio. (They consider the target debt to be 1/2 the credit line or less). Jan Feb March April

HelpMePlz:

Everytime I try to figure it out, it never gets paid off cause the interest just makes it go up and up

Champion:

There may exist a formula for this problem, but it might be simpler just do the month to month calculations, since the maximum number of months is four, and minimum 2. Target debt is $900/2=450 or less. ASSUME payment is at the end of the month, then At the end of January: Debt carried over=$498 add interest = $498*0.15/12)=$6.30; total debt=498+6.30=504.30 debt after payment = 504.30-35=$469.22 At the end of February: .... continue as above until debt after payment is at or below $450.

HelpMePlz:

ooooh i was doing 15% every month

HelpMePlz:

thank you!

Champion:

Interest rate of 15$ (0.15) applies to all year, yes, for one month it is 0.15/12!

Champion:

you're welcome ^_^